OpenStudy (anonymous):

Find the height of a cone with volume 4034.66 cm3 and radius 13 cm.

OpenStudy (anonymous):

volume of cone=1/3 pi*r^2*h

OpenStudy (anonymous):

4034.66=1/3*pi*13*13*h

OpenStudy (anonymous):

find h now

OpenStudy (anonymous):

I got h = 2.53. However, that isn't one of my answer choices. Here are the choices I have: 4.56 cm 11.4 cm 22.8 cm 34.2 cm

OpenStudy (anonymous):

nope .check it again

OpenStudy (anonymous):

4034.66=(1/3)*pi*13*13*h

OpenStudy (anonymous):

Oh! I was dividing by 3 instead of multiplying by 3. I got 22.8 cm.

OpenStudy (anonymous):

yes.u got it right now

OpenStudy (anonymous):

pi=3.14 or 22/7